OpenStudy (vera_ewing):

What is the y-value of tan(x) when x = -180°?

OpenStudy (vera_ewing):

@zzr0ck3r would the answer be -1?

OpenStudy (zzr0ck3r):

\(\large \tan(x) = \frac{\sin(x)}{\cos(x)}\)

OpenStudy (zzr0ck3r):

what is \(\sin(-180)\)?

OpenStudy (vera_ewing):

i'm not sure

OpenStudy (zzr0ck3r):

you just answered this question

OpenStudy (vera_ewing):

but i'm not really sure. i just guessed.

OpenStudy (zzr0ck3r):

then you need to look at a unit circle.

OpenStudy (vera_ewing):

ok can you draw one?

OpenStudy (zzr0ck3r):

these are the very basics of trig, and you should not move on until you get them, but once you do it will make the reast easy

OpenStudy (vera_ewing):

yeah i'm a little worried because i don't understand the basics

OpenStudy (zzr0ck3r):

http://www.youtube.com/watch?v=ZffZvSH285c

OpenStudy (zzr0ck3r):

watch that.

OpenStudy (vera_ewing):

ok thanks!

OpenStudy (zzr0ck3r):

this guy is great. world famous.

OpenStudy (vera_ewing):

good, thanks. this should help :)

OpenStudy (zzr0ck3r):

get back to me when you are done